The volume of the cube is 8 cubic cm. If you double the length, width and height of the cube, what would the new volume be?
Answer by mananth(16949)

the present volume is 8 cu. cm
volume = 2*2*2 =8
side =2 cm
increase by 2 cm
side will be 4 cm
4*4*4 = 64 cu. cm will be new volume